OpenStudy (anonymous):

h(t)=-16t^2+voT+ho. Initial velocity is 64ft/s. The rocket hits the ground in 4 seconds. If the rocket reached its maximum height (vertex) t=2 seconds, what is the value of maximum height?

hero (hero):

I'm pretty sure the initial height is zero, so you don't even count it

hero (hero):

4 = -16t^2+64t Now we just solve for t

hero (hero):

-16t^2+64t - 4 = 0

OpenStudy (anonymous):

-32t^2-256t-16

OpenStudy (anonymous):

never mind..im suppose to factor right?

hero (hero):

-4(4t^2-16t+1) = 0 4t^2-16+1 = 0

OpenStudy (anonymous):

is it 96 feet?

hero (hero):

4t^2+(-8 + 2sqrt{15})x + (-8 - 2sqrt{15})x + 1 = 0

OpenStudy (anonymous):

confused...

OpenStudy (anonymous):

ok

hero (hero):

Belay that

OpenStudy (anonymous):

Can you take a look at this Hero when you get a chance. Thanks http://openstudy.com/?F47433677217634HJ3W=_#/updates/4eb9de7ce4b0fc85a8ca99ac

OpenStudy (anonymous):

belay?

OpenStudy (anonymous):

okay I got it..thanks

OpenStudy (anonymous):

tht was a complicated problem..thanx for helping